OpenStudy (anonymous):

- tan ^ 2 x+ sec ^2 x = 1

OpenStudy (jdoe0001):

\(\bf sec^2(x)-tan^2(x) \implies \cfrac{1}{cos^2(x)}-\cfrac{sin^2(x)}{cos^2(x)} \implies \cfrac{1-sin^2(x)}{cos^2(x)}\) check your trig identities to see what \(\bf 1-sin^2(x)\) is

OpenStudy (anonymous):

Its negative tan though

OpenStudy (jdoe0001):

\(\bf -tan^2(x)+sec^2(x)=1 \implies sec^2(x)-tan^2(x)\)

OpenStudy (anonymous):

Ok so it flips places

OpenStudy (jdoe0001):

well, yes, I used that, but yes, I commute it :)

OpenStudy (anonymous):

Ok would 1-sin ^2 (x) = cos x

OpenStudy (jdoe0001):

well \(\bf sin^2(\theta)+cos^2(\theta)=1\) if you solve for \(\bf cos^2(\theta)\) what do you get?

OpenStudy (anonymous):

I don't know

OpenStudy (jdoe0001):

how about this one a + b = 1 if you solve for "b", what do you get?

OpenStudy (anonymous):

0

OpenStudy (jdoe0001):

0?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

u still here

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

Ok well could u still guide me through this ?

OpenStudy (jdoe0001):

well, there's no much there really, you see \(\bf sin^2(\theta)+cos^2(\theta)=1\) then see what you get if you were to solve for \(\bf cos^2(\theta)\)

OpenStudy (anonymous):

What do you mean ? Would I subtract cos from both sides? I am brain farting right now

OpenStudy (jdoe0001):

well, to isolate \(\bf cos^2(\theta)\) you'd need to substract \(\bf sin^2(\theta)\) from both sides

OpenStudy (anonymous):

Ok wow so would \[\cos ^{2}(\theta) = 1 - \sin ^{2} (\theta)\] be the answer